40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Nov 10 12:41:23 2020, gautam, Update, IOO, 1W > IMC CD93A725-FB5C-4F67-BB2E-D122205114B0.jpeg
    Reply  Tue Nov 10 14:30:06 2020, gautam, Update, IOO, WFS2 broken  WFS2broken.pngWFS2broken_RF.png
       Reply  Fri Dec 4 15:58:22 2020, Koji, Update, IOO, WFS2 broken  
Message ID: 15669     Entry time: Tue Nov 10 12:41:23 2020     Reply to this: 15670
Author: gautam 
Type: Update 
Category: IOO 
Subject: 1W > IMC 

Looking back through the elog, 1mtorr is the pressure at which it is deemed safe to send the full power beam into the IMC. After replacing the HR mirror in the MCREFL path with a 10% reflective BS, I just cranked the power back up. IMC is locked. With the increased exposure on the MC2T camera, lots of new scattered light has become visible.

Attachment 1: CD93A725-FB5C-4F67-BB2E-D122205114B0.jpeg  3.499 MB  | Hide | Hide all
CD93A725-FB5C-4F67-BB2E-D122205114B0.jpeg
ELOG V3.1.3-